OpenStudy (anonymous):

help me pleaseeeeeeeeeeeeeeeeeeeeee.. if sin A + Sin^2 A = 1, find the value of: (1) sin A (2) cos^2 A + 2 cos^4 A.

OpenStudy (anonymous):

1) SINA( 1+ SINA) =1 SO SINA = 1 SINA=0 A= 90 A= 180 ,0

OpenStudy (anonymous):

i did that.. that is not the answer that they want..i have the answer actually.. the answer for (1) is \[(-1+\sqrt{5}) \div 2\]

OpenStudy (anonymous):

i have all the answer but i don't know how to do..

OpenStudy (anonymous):

Use Quadratic equation here..

OpenStudy (anonymous):

\[\sin^2(A) + \sin(A) - 1 = 0\] a = 1, b= 1 and c = -1 Find Discriminant here: \[D = b^2 - 4ac = 1 + 4 = 5\] \[\sqrt{D} = \sqrt{5}\] \[\sin(A) = \frac{-b \pm \sqrt{D}}{2a}\] \[\sin(A) = \frac{-1 \pm \sqrt{5}}{2}\]

OpenStudy (anonymous):

Getting @SyaIE

OpenStudy (anonymous):

ohhhhhhhhhhhhhhhhhhh!! i get it!! thank you~ :)

OpenStudy (anonymous):

Can you tell me answer of your Second Part..

OpenStudy (anonymous):

YES YOU ARE RIGHT ITHINK YOU WANT FIND A SORY

OpenStudy (anonymous):

@obada its okay.. thanks anyway.. answer for (2) is \[(5+\sqrt{5}) \div 2\]

OpenStudy (anonymous):

Are you able to do second part ??

OpenStudy (anonymous):

ermm.. not really.. im still doing it..

OpenStudy (anonymous):

Think like this: \[\cos^2(A) + 2 \cos^4(A) \implies \cos^2(A)(1 + 2\cos^2(A))\] I am still thinking how to do easily.. So give me sometime..

OpenStudy (anonymous):

what i did was\[1-\cos ^{2} A + (1-\cos ^{2})^{2} =1\] then i expand the \[(1-\cos ^{2} A)^{2}\].. now im stuck

OpenStudy (anonymous):

Hey you have solved it.. I think you did not get how I tell you..

OpenStudy (anonymous):

yeah.. i don't get it, i guess.. hihi

OpenStudy (anonymous):

@waterineyes gooooood teacher

OpenStudy (anonymous):

See: \[(1 - \sin^2(A)) + 2(\cos^2(A))^2 \implies (1 - \sin^2(A)) + 2(1 - \sin^2(A))^2\] \[\implies (1- \sin^2(A))(1 +2(1-\sin^2(A))) \implies (1 - \sin^2(A)) \times (3 - 2\sin^2(A))\]

OpenStudy (anonymous):

Give me some more time.. Now I am also stuck..

OpenStudy (anonymous):

We have to convert that cos terms into sin so that we can use the given.. So let me try first..

OpenStudy (anonymous):

okie~:)

OpenStudy (anonymous):

I am close but I am trying now also to get the right answer..

OpenStudy (anonymous):

i think im about get the answer.

OpenStudy (anonymous):

See we are given with: \[\sin(A) + \sin^2(A) = 1\] So: \[\sin(A) = 1 - \sin^2(A) = \cos^2(A)\] Till here you got??

OpenStudy (anonymous):

\[1 - \sin^2(A) = \cos^2(A)\] This is the basic identity.. Getting??

OpenStudy (anonymous):

So we have: \[\sin(A) = \cos^2(A)\]

OpenStudy (anonymous):

yup..

OpenStudy (anonymous):

\[\cos^2(A) + 2(\cos^2(A))^2 \implies \sin(A) + 2\sin^2(A)\] Getting till here..

OpenStudy (anonymous):

Having any problem till here?/ @SyaIE

OpenStudy (anonymous):

no problem..i get it.. :)

OpenStudy (anonymous):

Can you go forward now on your own??

OpenStudy (anonymous):

You can write it as: \[\sin(A) + 2\sin^2(A) = \color{green}{(\sin(A) + \sin^2(A))} + \sin^2(A)\] \[\implies 1 + \sin^2(A)\]

OpenStudy (anonymous):

alright...

OpenStudy (anonymous):

So to find \(sin^2(A)\) You have sin(A) = : \[\sin(A) = \frac{-1 + \sqrt{5}}{2}\] Square it: \[\sin^2(A) = \frac{1 + 5 + 2 \sqrt{5}}{4} \implies \frac{3+ \sqrt{5}}{2}\] Add it to 1 now and you will have your answer..

OpenStudy (anonymous):

yupp!!.. i get it.. but.. i got.\[(3-\sqrt{5})\div2\]

OpenStudy (anonymous):

Or you can have this also: \[\sin^2(A) = \frac{1 + 5 - 2 \sqrt{5}}{4} \implies \frac{3- \sqrt{5}}{2}\]

OpenStudy (anonymous):

Well Done..

OpenStudy (anonymous):

See: \[1 + \frac{3 - \sqrt{5}}{2} \implies \frac{2 + 3 - \sqrt{5} }{2} \implies \frac{5 - \sqrt{5}}{2}\]

OpenStudy (anonymous):

yeah.. i get it~.. THANK YOU SOOOOOOOOOOOOOOOOOOOO MUCH..:)

OpenStudy (anonymous):

So your answer will be both: \[\frac{5 \pm \sqrt{5}}{2}\]

OpenStudy (anonymous):

YUP..

OpenStudy (anonymous):

Welcome dear..